Three point charges are located on the positive x-axis of a coordinate system. Charge q1 = 2.0 nC is 2.0 cm from the origin, charge q2 = -4.0 nC is 4.0 cm from the origin and charge q3 = 4.0 nC located at the origin.

A. What is the magnitude of the net force exerted by the other two charges on charge q1 = 2.0 nC?

Express your answer in newtons.

B.What is the direction of the net force exerted by the other two charges on charge q1 = 2.0 nC?

Express your answer in degrees.
Answer #1

here,

q1 = 2 nC = 2 * 10^-9 C is at x1 = 2 cm = 0.02 m

q2 = -4 nC = -4 * 10^-9 C is at x2 = 4 cm = 0.04 m

q3 = 4 nC = 4 * 10^-9 C is at x3 = 0 m

a)

the magnitude of the net force exerted by the other two charges on q1 , F = F12 + F13

F = K * q1 * q2 /(x2 - x1)^2 + K * q1 * q3 /x1^2

F = 9 * 10^9 * 2 * 10^-9 * 10^-9 * ( 4 /0.02^2 + 4 /0.02^2) N

F = 3.6 * 10^4 N

b)

the direction of the net force is 0 degree from +X axis